Can I please have some help with the following
3x+9y=21
-8x+y=19

7m+n=1
m-2n=13

Answer by mananth(16949)

3x+9y=21 .............1
-8x+y=19 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -9
3x+9y=21
72x-9y=-171
Add the two equations
75x= -150
/75
x=-2
plug value of x in (1)
3x+9y=21
-6+9y=21
9y=21+6
9y=27
y=3
...................

7m+n=1 .............1
m-2n=13 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
14m+2n= 2
m-2n= 13
Add the two equations
15m= 15
/15
m= 1
plug value of m in (1)
7m+n=1
7+1n=1
n=1-7
n=-6